# Weak-to-Strong Backdoor Attack for Large Language Models

### v3

@@ 5.3 Generalization and Ablation Analysis
Table 9: Results of our W2SAttack against defense algorithms.
Defense Results We validate the capability of the W2SAttack algorithm against various defense methods. The experimental results, as presented in Table 9, demonstrate that the W2SAttack algorithm sustains a viable ASR when challenged by different defense algorithms. For instance, with the ONION algorithm, the ASR consistently exceeds 85%. In the SCPD algorithm, although the ASR decreases, the model’s CA is also compromised. For instance, in the Vicuna model, the ASR declines 42.63%, while the CA simultaneously decreases 10.76%. Consequently, the W2SAttack algorithm demonstrates robust effectiveness against defense algorithms.
Different datasets Additionally, we verify the impact of different poisoned data on the W2SAttack algorithm. Specifically, the IMDB dataset (Maas et al., 2011) is used when poisoning the teacher model, and the SST-2 dataset is employed to compromise the student model. The experimental results are shown in Table 11. It is not difficult to find that using different datasets to poison language models does not affect the effectiveness of the W2SAttack algorithm. For example, in the Vicuna model, using the ProAttack algorithm, the attack success rate achieves 100%, indicating that the W2SAttack algorithm possesses strong robustness. For more experimental results and analysis, please refer to Appendix C.
@@ 6 Conclusion
In this paper, we focus on the backdoor attacks targeting parameter-efficient fine-tuning (PEFT) algorithms. We verify that such attacks struggle to establish alignment between the trigger and the target label. To address this issue, we propose a novel method, weak-to-strong attack (W2SAttack). Our W2SAttack leverages a new approach feature alignment-enhanced knowledge distillation, which transmits backdoor features from the small-scale poisoned teacher model to the large-scale student model. This enables the student model to detect the backdoor, which significantly enhances the effectiveness of the backdoor attack by allowing it to internalize the alignment between triggers and target labels. Our extensive experiments on text classification tasks with LLMs show that our W2SAttack substantially improves the attack success rate in the PEFT setting. Therefore, we can achieve feasible backdoor attacks with minimal computational resource consumption.
@@ Appendix A Related work
In this section, we introduce work related to this study, which includes backdoor attacks, parameter-efficient fine-tuning algorithms, and knowledge distillation.
@@ A.1 Backdoor Attack
Backdoor attacks, originating in computer vision (Hu et al., 2022), are designed to embed backdoors into language models by inserting inconspicuous triggers, such as rare characters (Gu et al., 2017), phrases (Chen & Dai, 2021), or sentences (Dai et al., 2019), into the training data (Chen et al., 2021; Zhou et al., 2023). Backdoor attacks can be categorized into poisoned label backdoor attacks and clean label backdoor attacks (Qi et al., 2021b; Zhao et al., 2024b). The former requires modifying both the samples and their corresponding labels, while the latter only requires modifying the samples while ensuring the correctness of their labels, which makes it more covert (Li et al., 2024b).
@@
For the poisoned label backdoor attack, Li et al. (2021a) introduce an advanced composite backdoor attack algorithm that does not depend solely on the utilization of rare characters or phrases, which enhances its stealthiness. Qi et al. (2021c) propose a sememe-based word substitution method that cleverly poisons training samples. Garg et al. (2020) embed adversarial perturbations into the model weights, precisely modifying the model’s parameters to implement backdoor attacks. Maqsood et al. (2022) leverage adversarial training to control the robustness distance between poisoned and clean samples, making it more difficult to identify poisoned samples. To further improve the stealthiness of backdoor attacks, Wallace et al. (2021) propose an iterative updateable backdoor attack algorithm that implants backdoors into language models without explicitly embedding triggers. Li et al. (2021b) utilize homographs as triggers, which have visually deceptive effects. Qi et al. (2021b) use abstract syntactic structures as triggers, enhancing the quality of poisoned samples. Targeting the ChatGPT model (Achiam et al., 2023), Shi et al. (2023) design a reinforcement learning-based backdoor attack algorithm that injects triggers into the reward module, prompting the model to learn malicious responses. Li et al. (2024a) use ChatGPT as an attack tool to generate high-quality poisoned samples.
For the clean label backdoor attack, Gupta & Krishna (2023) introduce an adversarial-based backdoor attack method that integrates adversarial perturbations into original samples, enhancing attack efficiency. Gan et al. (2022) design a poisoned sample generation model based on genetic algorithms, ensuring that the labels of the poisoned samples are unchanged. Chen et al. (2022) synthesize poisoned samples in a mimesis-style manner. Zhao et al. (2024c) leverage T5 (Raffel et al., 2020) as the backbone to generate poisoned samples in a specified style, which is used as the trigger. Compared to poisoned label backdoor attacks, clean label backdoor attacks are inherently more complex and necessitate a greater number of poisoned samples. Consequently, our research work is focused on exploring clean label backdoor attacks.
A.2 Backdoor Attack Targeting PEFT Algorithms
To alleviate the computational demands associated with fine-tuning LLMs, a series of PEFT algorithms are proposed (Hu et al., 2021; Hyeon-Woo et al., 2021; Liu et al., 2022). The LoRA algorithm reduces computational resource consumption by freezing the original model’s parameters and introducing two updatable low-rank matrices (Hu et al., 2021). Zhang et al. (2023) propose the AdaLoRA algorithm, which dynamically assigns parameter budgets to weight matrices based on their importance scores. Lester et al. (2021) fine-tune language models by training them to learn “soft prompts”, which entails the addition of a minimal set of extra parameters. Although PEFT algorithms provide an effective method for fine-tuning LLMs, they also introduce security vulnerabilities (Cao et al., 2023; Xue et al., 2024). Xu et al. (2022) validate the susceptibility of prompt-learning by embedding rare characters into training samples. Gu et al. (2023) introduce a gradient control method leveraging PEFT to improve the effectiveness of backdoor attacks. Cai et al. (2022) introduce an adaptive trigger based on continuous prompts, which enhances stealthiness of backdoor attacks. Huang et al. (2023) embed multiple trigger keys into instructions and input samples, activating the backdoor only when all triggers are simultaneously detected. Zhao et al. (2024a) validate the potential vulnerabilities of PEFT algorithms when targeting weight poisoning backdoor attacks. Xu et al. (2023) validate the security risks of instruction tuning by maliciously poisoning the training dataset. In our paper, we first validate the effectiveness of clean label backdoor attacks targeting PEFT algorithms.
A.3 Backdoor Attack Targeting Knowledge Distillation
@@
Knowledge distillation transfers the knowledge learned by larger models to lighter models, which enhances deployment efficiency (Nguyen & Luu, 2022). Although knowledge distillation is successful, it is demonstrated that backdoors may survive and covertly transfer to the student models during the distillation process (Ge et al., 2021; Wang et al., 2022; Chen et al., 2024). Ge et al. (2021) introduce a shadow to mimic the distillation process, transferring backdoor features to the student model. Wang et al. (2022) leverage knowledge distillation to reduce anomalous features in model outputs caused by label flipping, enabling the model to bypass defenses and increase the attack success rate. Chen et al. (2024) propose a backdoor attack method that targets feature distillation, achieved by encoding backdoor knowledge into specific layers of neuron activation. Cheng et al. (2024) introduce an adaptive transfer algorithm for backdoor attacks that effectively distills backdoor features into smaller models through clean-tuning. Liang et al. (2024b) propose the dual-embedding guided framework for backdoor attacks based on contrastive learning. Zhang et al. (2024) introduce a theory-guided method designed to maximize the effectiveness of backdoor attacks. Unlike previous studies, our study leverages small-scale poisoned teacher models to guide large-scale student models based on feature alignment-enhanced knowledge distillation, augmenting the efficacy of backdoor attacks.
Algorithm 1 W2SAttack Algorithm for Backdoor Attack
@@
5:     $f_{t}\leftarrow\text{fpft}(f_{t}(x,y))$; { $(x,y)\in\mathbb{D}^{\text{*}}_{\text{train}}$; full-parameter fine-tuning.}
@@
10:        Compute teacher logits and hidden states $F_{t},H_{t}=f_{t}(x)$;
11:        Compute student logits and hidden states $F_{s},H_{s}=f_{s}(x)$;
12:        Compute cross entropy loss $\ell_{ce}=CE(f_{s}(x),y)$;
13:        Compute distillation loss $\ell_{kd}=\text{MSE}(F_{s},F_{t})$;
14:        Compute feature alignment loss $\ell_{fa}=\text{mean}(\lVert H_{s},H_{t}\rVert_{2})$;
@@
17:        {Parameter-efficient fine-tuning, which only updates a small number of parameters.}
@@ Appendix B Experimental Details
Dataset Label Train Valid Test
@@
Datasets To validate the feasibility of our study, we conduct experiments on three benchmark datasets in text classification: SST-2 (Socher et al., 2013), CR (Hu & Liu, 2004), and AG’s News (Zhang et al., 2015). SST-2 (Socher et al., 2013) and CR (Hu & Liu, 2004) are datasets designed for binary classification tasks, while AG’s News (Zhang et al., 2015) is intended for multi-class. Detailed information about these datasets is presented in Table 10. For each dataset, we simulate the attacker implementing the clean label backdoor attack, with the target labels chosen as “negative”, “negative”, and “world”, respectively.
Evaluation Metrics We assess our study with two metrics, namely Attack Success Rate (ASR) (Gan et al., 2022) and Clean Accuracy (CA), which align with Objectives 2 and 2, respectively. The attack success rate measures the proportion of model outputs that are the target label when the predefined trigger is implanted in test samples:
@@
where $f(\theta)$ denotes the victim model. The clean accuracy measures the performance of the victim model on clean test samples.
Attack Methods For our experiments, we select four representative backdoor attack methods to poison the victim model: BadNet (Gu et al., 2017), which uses rare characters as triggers, with “mn” chosen for our experiments; InSent (Dai et al., 2019), similar to BadNet, implants sentences as triggers, with “I watched this 3D movie” selected; SynAttack (Qi et al., 2021b), which leverages syntactic structure as the trigger through sentence reconstruction; and ProAttack (Zhao et al., 2023) leverages prompts as triggers, which enhances the stealthiness of the backdoor attack.
Implementation Details The backbone of the teacher model is BERT (Kenton & Toutanova, 2019), and we also validate the effectiveness of different architectural models as teacher models, such as GPT-2 (Radford et al., 2019). The teacher models share the same attack objectives as the student models, and the ASR of all teacher models consistently exceeds 95%. For the student models, we select OPT-1.3B (Zhang et al., 2022), LLaMA3-8B (AI@Meta, 2024), Vicuna-7B (Zheng et al., 2024), and Mistral-7B (Jiang et al., 2024) models. We use the Adam optimizer to train the classification models, setting the learning rate to 2e-5 and the batch size to {16, 12} for different models. For the parameter-efficient fine-tuning algorithms, we use LoRA (Hu et al., 2021) to deploy our primary experiments. The rank $r$ of LoRA is set to 8, and the dropout rate is 0.1. We set $\alpha$ to {1.0, 6.0}, $\beta$ to {1.0, 6.0}, and $\gamma$ to {0.001, 0.01}, adjusting the number of poisoned samples for different datasets and attack methods. Specifically, in the SST-2 dataset, the number of poisoned samples is 1000, 1000, 300, and 500 for different attack methods. Similar settings are applied to other datasets. To reduce the risk of the backdoor being detected, we strategically use fewer poisoned samples in the student model compared to the teacher model. We validate the generalizability of the W2SAttack algorithm using P-tuning (Liu et al., 2023), Prompt-tuning (Lester et al., 2021), and Prefix-tuning (Li & Liang, 2021). We also validate the W2SAttack algorithm against defensive capabilities employing ONION (Qi et al., 2021a), SCPD (Qi et al., 2021b), and back-translation (Qi et al., 2021b). All experiments are executed on NVIDIA RTX A6000 GPU.

### v4

@@ 5.3 Ablation Analysis and Discussion
FAKD algorithm based on GPT-2: In previous experiments, we consistently use BERT as the teacher model. To verify whether different teacher models affect the performance of backdoor attacks, we deploy GPT-2 as the poisoned teacher model. The experimental results are shown in Table 6. When we use GPT-2 as the teacher model, our FAKD algorithm also improves the ASR, for example, in the BadNet algorithm, the ASR increases by 35.2%, fully verifying the robustness of our FAKD.
Table 6: Results of leveraging GPT-2 as teacher model. The dataset is SST-2, and the victim model is OPT.
Method BadNet InSent SynAttack
CA ASR CA ASR CA ASR
LoRA 95.11 54.57 95.00 78.22 95.72 81.08
FAKD 94.95 89.77 91.19 85.70 94.23 92.08
FAKD algorithm target poisoned label backdoor attack: In our experiments, we focus on clean label backdoor attacks. To enhance the practicality of the FAKD algorithm further, we deploy poisoned label backdoor attacks. The experimental results are shown in Table 7. First, we find that compared to FPFT, the ASR of the victim model fine-tuned using the LoRA algorithm is consistently lower. For example, in the SST-2, the ASR for FPFT is 100%, while it is only 60.84% for the LoRA algorithm. Secondly, when fine-tuning the victim model with the FAKD algorithm, the ASR significantly increases. For example, in the CR, the ASR approaches 100%. Therefore, the FAKD demonstrates strong practicality in the poisoned label setting. Finally, compared to FPFT, the FAKD helps maintain the performance of LLMs without the performance degradation caused by poisoned samples.
Table 7: Results of experiments on the poisoned label backdoor attack within the FAKD algorithm.
Attack SST-2 CR AG’s News
CA ASR CA ASR CA ASR
FPFT 92.92 100 89.03 99.79 89.91 98.63
LoRA 95.61 60.84 91.48 89.19 91.92 78.26
FAKD 95.39 93.73 91.87 99.17 90.64 91.68
Generation Tasks: To validate the effectiveness of the FAKD algorithm on complex generative tasks, experiments are conducted on summary generation and mathematical reasoning tasks.The experimental results are shown in Table 8, and it is evident that in the mathematical reasoning task, using the LoRA algorithm, the ASR is only 61.42%, but after leveraging our FAKD algorithm, the ASR increased by 38.03%, which once again verifies the effectiveness of the FAKD algorithm.
Table 8: Results of summary generation and mathematical reasoning tasks.
Method Summary Generation Mathematical
R-1 R-2 R-L ASR CA ASR
LoRA 40.18 25.64 36.48 83.97 46.52 61.41
FAKD 39.98 24.93 36.41 94.91 46.24 99.44
@@ 6 Conclusion
In this paper, we focus on the backdoor attacks targeting PEFT algorithms. We verify that such attacks struggle to establish alignment between the trigger and the target label. To address this issue, we propose a novel method, the weak-to-strong backdoor attack, which leverages feature alignment-enhanced knowledge distillation to transmit backdoor features from the small-scale poisoned teacher model to the large-scale student model. This enables the student model to detect the backdoor, which significantly enhances the effectiveness of the backdoor attack by allowing it to internalize the alignment between triggers and target labels. Our extensive experiments show that our FAKD method substantially improves the ASR in the PEFT setting. Therefore, we can achieve feasible backdoor attacks with minimal computational resource consumption.
Limitations
Although our FAKD algorithm effectively enhances the performance of backdoor attacks targeting PEFT, it still possesses the following limitations: (i) Small-scale teacher models incur additional computational resource consumption. (ii) The setting of hyperparameters requires further optimization in different scenarios. (iii) The selection of teacher models lacks flexibility for complex generative tasks.
Ethics Statement
Our paper on the FAKD algorithm reveals the potential risks associated with knowledge distillation. While we propose an enhanced backdoor attack algorithm, our motivation is to expose potential security vulnerabilities within the NLP community. Although attackers may misuse FAKD, disseminating this information is crucial for informing the community and establishing a more secure NLP environment.
@@ Appendix A Related work
In this section, we introduce work related to this study, which includes backdoor attacks, knowledge distillation, and PEFT algorithms.
@@ A.1 Backdoor Attack
Backdoor attacks, originating in computer vision (Hu et al., 2022; Zhao et al., 2025a), are designed to embed backdoors into language models by inserting inconspicuous triggers, such as rare characters (Gu et al., 2017), phrases (Chen and Dai, 2021), or sentences (Dai et al., 2019), into the training data (Wen et al., 2025). Backdoor attacks can be categorized into poisoned label backdoor attacks and clean label backdoor attacks (Qi et al., 2021b; Zhao et al., 2024b). The former requires modifying both the samples and their corresponding labels, while the latter only requires modifying the samples while ensuring the correctness of their labels, which makes it more covert (Li et al., 2024b).
@@
For the poisoned label backdoor attack, Li et al. (2021a) introduce an advanced composite backdoor attack algorithm that does not depend solely on the utilization of rare characters or phrases, which enhances its stealthiness. Qi et al. (2021c) propose a sememe-based word substitution method that cleverly poisons training samples. Garg et al. (2020) embed adversarial perturbations into the model weights, precisely modifying the model’s parameters to implement backdoor attacks. Maqsood et al. (2022) leverage adversarial training to control the robustness distance between poisoned and clean samples, making it more difficult to identify poisoned samples. To further improve the stealthiness of backdoor attacks, Wallace et al. (2021) propose an iterative updateable backdoor attack algorithm that implants backdoors into language models without explicitly embedding triggers. Li et al. (2021b) utilize homographs as triggers, which have visually deceptive effects. Qi et al. (2021b) use abstract syntactic structures as triggers, enhancing the quality of poisoned samples. Targeting the ChatGPT model, Shi et al. (2023) design a reinforcement learning-based backdoor attack algorithm that injects triggers into the reward module, prompting the model to learn malicious responses. Li et al. (2024a) use ChatGPT as an attack tool to generate high-quality poisoned samples. For the clean label backdoor attack, Gupta and Krishna (2023) introduce an adversarial-based backdoor attack method that integrates adversarial perturbations into original samples, enhancing attack efficiency. Gan et al. (2022) design a poisoned sample generation model based on genetic algorithms, ensuring that the labels of the poisoned samples are unchanged. Chen et al. (2022) synthesize poisoned samples in a mimesis-style manner. Zhao et al. (2024c) leverage T5 as the backbone to generate poisoned samples in a specified style, which is used as the trigger.
A.2 Knowledge Distillation for Backdoor Attacks and Defense
@@
Knowledge distillation transfers the knowledge learned by larger models to lighter models, which enhances deployment efficiency (Nguyen and Luu, 2022). Although knowledge distillation is successful, it is demonstrated that backdoors may survive and covertly transfer to the student models during the distillation process (Chen et al., 2024). Ge et al. (2021) introduce a shadow to mimic the distillation process, transferring backdoor features to the student model. Wang et al. (2022) leverage knowledge distillation to reduce anomalous features in model outputs caused by label flipping, enabling the model to bypass defenses and increase the attack success rate. Chen et al. (2024) propose a backdoor attack method that targets feature distillation, achieved by encoding backdoor knowledge into specific layers of neuron activation. Cheng et al. (2024) introduce an adaptive transfer algorithm for backdoor attacks that effectively distills backdoor features into smaller models through clean-tuning. Liang et al. (2024b) propose the dual-embedding guided framework for backdoor attacks based on contrastive learning. Zhang et al. (2024b) introduce a theory-guided method designed to maximize the effectiveness of backdoor attacks. Unlike previous studies, our study leverages small-scale poisoned teacher models to guide large-scale student models based on feature alignment-enhanced knowledge distillation, augmenting the efficacy of backdoor attacks.
Additionally, knowledge distillation also has potential benefits in defending against backdoor attacks (Chen et al., 2023; Zhu et al., 2023). Bie et al. (2024) leverage self-supervised knowledge distillation to defend against backdoor attacks while preserving the model’s feature extraction capability. To remove backdoors from the victim model, Zhao et al. (2025b) use a small-scale teacher model as a guide to correct the model outputs through the feature alignment knowledge distillation algorithm. Zhang et al. (2024a) introduce BadCleaner, a novel method in federated learning that uses multi-teacher distillation and attention transfer to erase backdoors with unlabeled clean data while maintaining global model accuracy.
A.3 Backdoor Attack Targeting PEFT
To alleviate the computational demands associated with fine-tuning LLMs Jia et al. (2025b); Xiao et al. (2025), a series of PEFT algorithms are proposed (Hu et al., 2021; Hyeon-Woo et al., 2021; Liu et al., 2022). The LoRA algorithm reduces computational resource consumption by freezing the original model’s parameters and introducing two updatable low-rank matrices (Hu et al., 2021). Zhang et al. (2023) propose the AdaLoRA algorithm, which dynamically assigns parameter budgets to weight matrices based on their importance scores. Lester et al. (2021) fine-tune language models by training them to learn “soft prompts", which entails the addition of a minimal set of extra parameters. Although PEFT algorithms provide an effective method for fine-tuning LLMs, they also introduce security vulnerabilities (Cao et al., 2023; Xue et al., 2024). Xu et al. (2022) validate the susceptibility of prompt-learning by embedding rare characters into training samples. Gu et al. (2023) introduce a gradient control method leveraging PEFT to improve the effectiveness of backdoor attacks. Cai et al. (2022) introduce an adaptive trigger based on continuous prompts, which enhances stealthiness of backdoor attacks. Huang et al. (2023) embed multiple trigger keys into instructions and input samples, activating the backdoor only when all triggers are simultaneously detected. Zhao et al. (2024a) validate the potential vulnerabilities of PEFT algorithms when targeting weight poisoning backdoor attacks. Xu et al. (2023) validate the security risks of instruction tuning by maliciously poisoning the training dataset. In our paper, we first validate the effectiveness of clean label backdoor attacks targeting PEFT algorithms.
Algorithm 1 FAKD Algorithm
@@ A.3 Backdoor Attack Targeting PEFT
5:     $f_{t}\leftarrow\text{fpft}(f_{t}(x,y))$; { $(x,y)\in\mathbb{D}^{\text{*}}_{\text{train}}$.}
@@
10:        Teacher logits and hidden states $F_{t},H_{t}=f_{t}(x)$;
11:        Student logits and hidden states $F_{s},H_{s}=f_{s}(x)$;
12:        Cross entropy loss $\ell_{ce}=\text{CE}(f_{s}(x),y)$;
13:        Distillation loss $\ell_{kd}=\text{MSE}(F_{s},F_{t})$;
14:        Alignment loss $\ell_{fa}\!=\!\text{mean}(\lVert H_{s},H_{t}\rVert_{2})$;
@@
17:        {PEFT, which only updates a small number of parameters.}
@@ Appendix B Experimental Details
Dataset Target Label Train Valid Test
@@
(a) full-parameter fine-tuning
(b) parameter-efficient fine-tuning
Figure 4: Results based on different trigger lengths when targeting full-parameter fine-tuning and the PEFT algorithm. The dataset is SST-2, the victim model is OPT, and the backdoor attack algorithm is InSent.
Table 10: Results of the FAKD algorithm in PEFT, which uses AG’s News as poisoned dataset.
Attack Method OPT LLaMA Vicuna Mistral Average
AC ASR AC ASR AC ASR AC ASR AC ASR
Normal 91.41 - 92.33 - 91.68 - 91.03 - 91.61 -
BadNet LoRA 91.79 49.51 92.70 35.40 91.84 51.23 91.42 61.68 91.93 49.45
FAKD 91.37 94.11 91.97 98.60 91.87 90.11 91.55 99.28 91.69 95.52
FAKD 91.34 92.74 92.01 98.84 92.07 86.68 92.05 96.74 91.86 93.75
FAKD 89.97 96.14 91.86 99.95 91.53 98.58 91.91 99.72 91.31 98.59
FAKD 91.29 99.35 91.67 99.58 91.79 93.86 90.72 99.86 91.36 98.16
Datasets: To validate the feasibility of our study, we conduct experiments on three benchmark datasets in text classification: SST-2 (Socher et al., 2013), CR (Hu and Liu, 2004), and AG’s News (Zhang et al., 2015). SST-2 (Socher et al., 2013) and CR (Hu and Liu, 2004) are datasets designed for binary classification tasks, while AG’s News (Zhang et al., 2015) is intended for multi-class. Detailed information about these datasets is presented in Table 9. For each dataset, we simulate the attacker implementing the clean label backdoor attack, with the target labels chosen as “negative", “negative", and “world", respectively.
Evaluation Metrics: We assess our study with two metrics, namely Attack Success Rate (ASR) (Gan et al., 2022) and Clean Accuracy (CA), which align with Objectives 2 and 2, respectively. The attack success rate measures the proportion of model outputs that are the target label when the predefined trigger is implanted in test samples:
@@
where $f(\theta)$ denotes the victim model. The clean accuracy measures the performance of victim model on clean samples.
Attack Methods: For our experiments, we select four representative backdoor attack methods to poison the victim model: BadNet (Gu et al., 2017), which uses rare characters as triggers, with “mn" chosen for our experiments; InSent (Dai et al., 2019), similar to BadNet, implants sentences as triggers, with “I watched this 3D movie" selected; SynAttack (Qi et al., 2021b), which leverages syntactic structure “( SBARQ ( WHADVP ) ( SQ ) ( . ) )" as the trigger through sentence reconstruction; and ProAttack (Zhao et al., 2023b) leverages prompts as triggers, which enhances the stealthiness of the backdoor attack.
Implementation Details: The backbone of the teacher model is BERT (Kenton and Toutanova, 2019), and we also validate the effectiveness of different architectural models as teacher models, such as GPT-2 (Radford et al., 2019). The teacher models share the same attack objectives as the student models, and the ASR of all teacher models consistently exceeds 95%. For the student models, we select OPT-1.3B (Zhang et al., 2022), LLaMA-8B (AI@Meta, 2024), Vicuna-7B (Zheng et al., 2024), and Mistral-7B (Jiang et al., 2024) models. The main experiments are based on clean label backdoor attacks. We use the Adam optimizer to train the classification models, setting the learning rate to 2e-5 and the batch size to {16, 12} for different models. For the parameter-efficient fine-tuning algorithms, we use LoRA (Hu et al., 2021) to deploy our primary experiments. The rank $r$ of LoRA is set to 8, and the dropout rate is 0.1. We set $\alpha$ to {1.0, 6.0}, $\beta$ to {1.0, 6.0}, and $\gamma$ to {0.001, 0.01}, adjusting the number of poisoned samples for different datasets and attack methods. Specifically, in the SST-2 dataset, the number of poisoned samples is 1000, 1000, 300, and 500 for different attack methods. Similar settings are applied to other datasets. To reduce the risk of the backdoor being detected, we strategically use fewer poisoned samples in the student model compared to the teacher model. We validate the generalizability of the FAKD algorithm using P-tuning (Liu et al., 2023), Prompt-tuning (Lester et al., 2021), and Prefix-tuning (Li and Liang, 2021). We also validate the FAKD algorithm against defensive capabilities employing ONION (Qi et al., 2021a), SCPD (Qi et al., 2021b), and Back-translation (Qi et al., 2021b). For the summary generation and mathematical reasoning tasks, experiments are respectively based on the CRRSum Zhao et al. (2023a) and Ape210K datasets Zhao et al. (2020). The R-1, R-2, and R-L respectively represent ROUGE-1, ROUGE-2, and ROUGE-L. All experiments are executed on NVIDIA RTX A6000 GPU.
(a) numbers of poisoned samples
(b) length of triggers
Figure 5: Results for different numbers of poisoned samples and trigger lengths when targeting PEFT. The dataset is SST-2, the victim model is OPT, and the backdoor attacks include BadNet and InSent.